OpenStudy (anonymous):

You deposit $6,700 in a money-market account that pays an annual interest rate of 3.3%.The interest is compounded quarterly. how much money will you have after 4 years?

OpenStudy (anonymous):

1st year 6700 + x1 ( x1 = (6700*3.3)/100 ) 2nd year 6700 + x1 + x2 ( x2 = ((6700+x1)*3.3)/00) and so on